# Modeling Domain and Feedback Transitions for Cross-Domain Sequential Recommendation

### v1

@@ Introduction
Table 2 and Table 3 demonstrate the performance of the methods in the “book-movie” and “book-music” CDSR scenarios.
@@
(2) For CDSR baselines, most baselines perform better than SR baselines because considering knowledge transfer between domains helps improve cross-domain recommendations. CoNet’s performance is slightly lower than SR baselines, possibly due to the weaker capability of its model architecture.
@@
Ablation Studies
Figure 3: Sensitivity Analysis of $L$.
Figure 4: Sensitivity Analysis of $N_{H}$.
The results of the ablation study, presented in Table 4 for the ”Book-Movie” scenario, reveal several key insights into the contributions of various components within $\text{Transition}^{2}$:
@@
  • Effect of Retaining Individual Transition Masks in Cross-Domain Sequence: When the model is configured to retain only individual transition masks in the MHSA of the cross-domain sequence (only $\mathbf{M}_{1}$, only $\mathbf{M}_{2}$, only $\mathbf{M}_{3}$, only $\mathbf{M}_{4}$), the general trend is a degradation in performance. This indicates that each transition mask contributes distinct information, and isolating them misses important cross-domain transitions. However, an interesting exception is observed with only $\mathbf{M}_{1}$ in the Movie domain, where a slight improvement is noted. This could imply that in certain cases, specific transitions might hold more relevance, but the overall performance gain from retaining all masks indicates that the comprehensive transition information is essential for robust recommendations.
@@
Hyper-parameters
@@
This section investigates the parameter sensitivity of the depth $L$ of the graph encoder and the number of self-attention heads $N_{H}$. For fair comparison, when studying $L$, we fix $N_{H}=8$; when studying $N_{H}$, we fix $L=1$. For the hyperparameter $L$, we conducted analysis experiments with values in {0, 1, 2, 3}, where $L=0$ means ignoring the Transition-Aware Graph Encoder. Figure 3 shows the impact on recommendation performance in terms of MRR@10 and NDCG@10 metrics for the “Book” and “Movie” domains, where it is observed that our model achieves the best performance when $L=1$ on Movie domain. For the hyperparameter $N_{H}$, we selected values from {4, 8, 12, 16}, and Figure 4 show the impact on recommendation performance in terms of MRR@10 and NDCG@10 metrics for the ”Book” and ”Movie” domains, where $\text{Transition}^{2}$ achieves the best recommendation performance on Book domain when $N_{H}=8$.
Conclusion
In this paper, we proposed $\text{Transition}^{2}$, a novel approach to model transitions across both domains and types of user feedback in cross-domain sequential recommendations. By using a transition-aware graph encoder and cross-transition multi-head self-attention, our method effectively captures and integrates the transition information from user history. Experimental results on two public datasets demonstrate the effectiveness of $\text{Transition}^{2}$ in improving prediction accuracy across different domains.

### v2

@@ 5.6. Overall Performance
Table 3, Table 4 and Table 5 demonstrate the performance of the methods in the “Book-movie” and “Book-music” of Douban, and “Book-Movie” of Amazon, both CDSR scenarios.
@@
(2) For CDSR baselines, most baselines perform better than SR baselines because considering knowledge transfer between domains helps improve cross-domain recommendations. CoNet’s performance is slightly lower than SR baselines, possibly due to the weaker capability of its model architecture. Moreover, TriCDR and CA-CDSR based on contrastive learning, exhibit superior performance. This indicates that contrastive learning holds great potential in enhancing representation learning and transferring knowledge within CDSR.
@@
Figure 3. Analysis of domain alignment loss $\mathcal{L}_{\text{align}}$.
Figure 4. Analysis of feedback contrast loss $\mathcal{L}_{\text{cont}}$.
5.7. Ablation Studies
The results of the ablation study, presented in Table 6 for the “Book-Movie” scenario, reveal several key insights into the contributions of various components within $\text{Transition}^{2}$:
@@ 5.7. Ablation Studies
  • Impact of Transition Alignment and Contrast: Removing intra-domain embedding alignment (w/o $\mathcal{L}_{\text{align}}$) or in-sequence positive-negative sample contrast (w/o $\mathcal{L}_{\text{cont}}$) leads to a significant decline in recommendation performance. This clearly demonstrates that embeddings within the same domain should exhibit stronger similarity, and positive/negative items from each user interaction should be differentiated to a greater extent. Both mechanisms contribute to improving the model’s performance by enhancing representation quality and discriminative power.
  • Effect of Retaining Individual Transition Masks in Cross-Domain Sequence: When the model is configured to retain only individual transition masks in the MHSA of the cross-domain sequence (only $\mathbf{M}_{1}$, only $\mathbf{M}_{2}$, only $\mathbf{M}_{3}$, only $\mathbf{M}_{4}$), the general trend is a degradation in performance. This indicates that each transition mask contributes distinct information, and isolating them misses important cross-domain transitions. However, an interesting exception is observed with only $\mathbf{M}_{1}$ in the Movie domain, where a slight improvement than others is noted. This could imply that in certain cases, specific transitions might hold more relevance, but the overall performance gain from retaining all masks indicates that the comprehensive transition information is essential for robust recommendations.
@@
5.8. Analysis of Transition Alignment and Contrast
5.8.1. Analysis of Domain Alignment Loss $\mathcal{L}_{\text{align}}$
To investigate the impact of the domain alignment loss $\mathcal{L}_{\text{align}}$ described in Section 4.4 on cross-domain representations, we first analyze its alignment effectiveness. As shown in Figure 3, we compare the cosine similarity distributions between Book domain embeddings $\mathbf{E}_{A}$ and Cross-domain embeddings $\mathbf{E}_{C}$, as well as between Movie domain embeddings $\mathbf{E}_{B}$ and $\mathbf{E}_{C}$, under conditions with and without $\mathcal{L}_{\text{align}}$. The significantly lower similarity observed without $\mathcal{L}_{\text{align}}$ demonstrates that our alignment loss effectively reduces embedding discrepancies caused by different encoders, thereby ensuring consistent cross-domain representation learning.
5.8.2. Analysis of Feedback Contrast Loss $\mathcal{L}_{\text{cont}}$
We further examine the effect of feedback contrast loss $\mathcal{L}_{\text{cont}}$ through inter-item similarity analysis. As illustrated in Figure 4, we calculate three similarity metrics using embeddings from Book domain: positive-positive similarity ($sim(p,p)$), positive-negative similarity ($sim(p,n)$), and negative-negative similarity ($sim(n,n)$). The results reveal that with $\mathcal{L}_{\text{cont}}$, same-feedback items exhibit enhanced intra-cluster similarity while showing reduced inter-cluster similarity (even reaching negative values in Movie domain). This contrastive pattern confirms that $\mathcal{L}_{\text{cont}}$ successfully differentiates user preference signals in interaction sequences, leading to more discriminative representations for recommendation improvement.
5.9. Sensitive Analysis of Hyper-parameters
Figure 5. Sensitivity Analysis of $L$.
Figure 6. Sensitivity Analysis of $N_{H}$.
In this section, we conducted more detailed sensitive analysis of the hyper-parameters setting in $\text{Transition}^{2}$.
5.9.1. Analysis of $L$ and $N_{H}$
@@
This section investigates the parameter sensitivity of the depth $L$ of the graph encoder and the number of self-attention heads $N_{H}$. For fair comparison, when studying $L$, we fix $N_{H}=8$; when studying $N_{H}$, we fix $L=1$. For the hyperparameter $L$, we conducted analysis experiments with values in {0, 1, 2, 3}, where $L=0$ means ignoring the Transition-Aware Graph Encoder. Figure 5 shows the impact on recommendation performance in terms of MRR@10 and NDCG@10 metrics for the “Book” and “Movie” domains on Douban, where it is observed that our model achieves the best performance when $L=1$ on Movie domain. For the hyperparameter $N_{H}$, we selected values from {4, 8, 12, 16}, and Figure 6 show the impact on recommendation performance in terms of MRR@10 and NDCG@10 metrics for the ”Book” and ”Movie” domains, where $\text{Transition}^{2}$  achieves the best recommendation performance on Book domain when $N_{H}=8$.
Figure 7. Sensitivity Analysis of $\tau$.
Figure 8. Sensitivity Analysis of $\mu_{1}$.
Figure 9. Sensitivity Analysis of $\mu_{2}$.
5.9.2. Analysis of $\tau$, $\mu_{1}$ and $\mu_{2}$
We analyze the impact of parameters $\tau$, $\mu_{1}$ and $\mu_{2}$ in the contrastive learning loss $\mathcal{L}_{\text{align}}$ and $\mathcal{L}_{\text{cont}}$. For fair comparison, when studying $\mu_{1}$ and $\mu_{2}$ with values in $\{0,0.2,0.5,0.7,1.0\}$, we fix $\tau=0.1$. $\mu_{1}=0$ and $\mu_{2}=0$ mean ignore the domain alignment loss $\mathcal{L}_{\text{align}}$ and the feedback contrast loss $\mathcal{L}_{\text{cont}}$, respectively; when studying $\tau$ with values in $\{0.01,0.05,0.1,0.5,1.0\}$, we fix $\mu_{1}=0.5$ and $\mu_{2}=0.5$. Figure 7, Figure 8, Figure 9 shows the impact of $\tau$, $\mu_{1}$, $\mu_{2}$ on recommendation performance in terms of MRR@10 and NDCG@10 metrics for the “Book” and “Movie” domains on Douban, respectively, where it can be observed from the figures that the optimal parameter settings on Book-Movie of Douban are observed to be $\mu_{1}=0.5$, $\mu_{2}=0.5$ and $\tau=0.1$. The results demonstrate that $\text{Transition}^{2}$ is quite sensitive to the hyper-parameters, and both too large and too small values of these parameters lead to a decrease in recommendation performance.
6. Conclusion
Our work addresses the limitations of current recommendation systems that predominantly focus on domain transitions by proposing $\text{Transition}^{2}$, a novel method to model user interest transitions across different domains and feedback types. By introducing a transition-aware graph encoder that assigns varying weights to edges based on feedback type, $\text{Transition}^{2}$  effectively captures the nuanced transition information in user history. The use of cross-transition multi-head self-attention further refines these embeddings by distinguishing different types of transitions with various masks. To further enhance representation learning, two contrastive losses are utilized to align transitions across domains and feedback types. Our experimental results on two public datasets validate the effectiveness of $\text{Transition}^{2}$, demonstrating its capability to enhance prediction accuracy across multiple domains. This approach not only advances the modeling of user interests in recommendation systems but also highlights the importance of incorporating both positive and negative feedback in cross-domain transitions.